OpenStudy (anonymous):

One x-intercept for a parabola is at the point (1,0). Use the factor method to find the other x-intercept for the parabola defined by this equation: y-5x^2+15x-10

OpenStudy (anonymous):

To find the x-intercept, set y = 0 and solve for x. Since the polynomial has an order of 2, and crosses the x-axis, it will have two roots. Find the roots of the equation using the quadratic equation, factoring, or complete the square.

myininaya (myininaya):

y=5(x^2+3x-2) 0=5(x^2+3x-2) 0=x^2+3x-2 (1,0) is not an x-intercept is your question written correctly?

myininaya (myininaya):

y=5x^2+15x-10 ?

OpenStudy (anonymous):

That's the exact question that's on the worksheet I have.